OpenStudy (anonymous):

Air in a piston/cylinder assembly at 200kpa and 600K is expanded in a constant-pressure process to twice the initial volume (state 2). The piston is then locked with a pin, and heat is transferred to a final temperature of 600K. (assume cp and cv are constant). A) find P, T and h for states 2 and 3 B) find 1W3 and 1Q3. I need help with B!!! i can't find volume to do the work equation, someone please help

OpenStudy (kainui):

If you know the information from A, can you use that along with PV=nRT to find the volume to help you in part B?

OpenStudy (anonymous):

My issue was that I don't know the value or n or V so I couldn't solve for the mass.

OpenStudy (kainui):

\[P_1V_1=nRT_1\]Let's just say this describes the first state. We're given some stuff already, \[P_1=200 \ kPa \\ T_1=600 \ K\] now they're saying that they expanded it to twice the original volume while keeping the pressure the same. That means since it's a closed container, the temperature had to change as well, so let's write it out \[P_2V_2=nRT_2\] But we can plug in a coupble values, since pressure was constant and volume doubled, we can write that mathematically as \[P_1=P_2 \\ V_2=2V_1\] Now we do a third thing too it by locking the pin and changing the temperature \[P_3V_3=nRT_3\] Locking the pin keeps the volume the same, and we changed the temperature back to the original temperature. So that means we can rewrite this as \[P_3V_2=nRT_1\] You can calculate the work without knowing the number of molecules since Work is just dependent on pressure and volume.

OpenStudy (anonymous):

so this is where i am at now: 1W3= 1W2+2W3 since volume is constant from process 2 to 3, 2W3=0 therefore: 1W3=1W2 1W2=(P1+P2)*(V2-V1)/2 I know the pressures, but i do not know the volumes Which equation do you think i can manipulate to find V2 or V1?

OpenStudy (anonymous):

YO welcome back Kai

OpenStudy (kainui):

Hmm this is interesting, I don't think there's enough information here to solve it or something lol. I thought I got the answer earlier but I think I was mistaken.

OpenStudy (kainui):

Can you leave it as a variable? Since V2=2*V1 then that means V2-V1 is simply V1 by itself. Hmm.

OpenStudy (anonymous):

While you are here, do you know how to solve this question? Suppose that X ∼ N(−7, 14). Find: (h) The value of x for which P(|X + 7| ≤ x) = 0.44